OpenStudy (zulema1030):

how many millilimeters in 23 inches?

OpenStudy (sshayer):

1 inch=2.54 cm=25.4 mm 23 inch=25.4 *23=?

OpenStudy (dopeness_):

584.2 Hope this help. c:

OpenStudy (mathmale):

Need to include the correct unit of measurement in your final answer.

OpenStudy (sanoofy):

23.4 in 1 inch. 25.4*23= 584.2

OpenStudy (zulema1030):

Thank you ❤️